Suppose Ford Motor Company issues bonds with a face value of $500 and an annual coupon payment of $5 What is the interest rate Ford is paying on the borrowed funds?
Answer #1

Coupon payment is the payment of annual interest to the bond holder. Here, the rate of interest = coupon payment/ face value * 100 = 5/500 * 100 = 1/100 * 100 = 1%